# AniMer+: Unified Pose and Shape Estimation Across Mammalia and Aves via Family-Aware Transformer

### v1

@@ V-B Qualitative Results of AniMer+
Figure 6: Qualitative comparisons on the Animal3D, CtrlAni3D, and Animal Kingdom datasets. We compare our results with HMR [19], WLDO [11], AniMer-a (ResNet152 backbone), AniMer-b (no pretraining) and HMR2.0 [15].
Figure 7: Qualitative comparisons on the CUB, CtrlAVES3D and Cow Bird datasets. We compare our results with AVES [2] and HMR2.0 [15].
@@ V-C1 Mammalian Experiments
Following the practice of the Animal3D dataset [18], we utilize HMR [19] and WLDO [11] as baselines. Also, we compare with HMR2.0 [15] to highlight our special design choices. To be fair, all compared methods are trained on our full training dataset with the same losses and the same two-stage training strategy.
@@
[…]
@@ V-C2 Avian Experiments
We conduct both quantitative and qualitative comparisons against AVES [2] and HMR2.0 [15]. Unlike mammals, birds exhibit a more limited range of poses and shapes, so the model converges faster. Therefore, we train AniMer-A in a single stage. The quantitative results in Table VI show that AniMer-A and AniMer+ outperform all baselines across all datasets. Notably, our method achieves the best results on the Cow Bird dataset despite its large number of blurred images and the fact that this dataset is not seen during training—underscoring the robustness of our method. While AniMer+ and AniMer-A perform comparably on our synthetic CtrlAVES3D dataset, the specialized AniMer-A achieves higher accuracy on the real-world CUB and Cow Bird datasets, as its entire network capacity is dedicated to avian reconstruction.
@@ V-E1 Ablations on Different Encoder and Decoder
To emphasize the significance of the ViT encoder, backbone pretraining and the Transformer decoder, we substitute the ViT encoder with a ResNet-152 encoder, discard ViT pretraining and replace the Transformer decoder with an MLP decoder, respectively. The results are presented in Table X. Compared with traditional methods, which use CNN as a backbone and an MLP head as a decoder, our ViT encoder and Transformer decoder improve the performance of the model’s output mesh and image alignment. In addition, a backbone with random initialization but no pretraining leads to poor performance and is hard to converge.
TABLE X: Ablations on different encoder and decoder. AniMer-a: use ResNet-152 as the encoder. AniMer-b: discard ViT pretraining. AniMer-c: use MLP as the decoder.
@@
PA-J $\downarrow$ PA-V $\downarrow$ PA-J $\downarrow$ PA-V $\downarrow$ AUC $\uparrow$ PCK@0.1 $\uparrow$
@@
V-E2 Ablation Study on Animal Family Supervised Contrastive Learning
TABLE XI: Effect of animal family supervised contrastive learning. PAJ: PA-MPJPE. PAV: PA-MPVPE. P@0.1: PCK@0.1.
@@ V-E2 Ablation Study on Animal Family Supervised Contrastive Learning
We conduct experiments to demonstrate the effect of animal family supervised contrastive learning. Table XI presents the quantitative results on both mammalian and avain datasets. Clearly, AniMer-M trained with $\mathcal{L}_{\text{con}}$ outperforms the model that does not utilize $\mathcal{L}_{\text{con}}$. The improved PA-MPVPE and PA-MPJPE metrics indicate that $\mathcal{L}_{\text{con}}$ enhances the model’s capability in shape and pose estimation.
@@
In addition, to provide more intuitive understanding of the effect of $\mathcal{L}_{\text{con}}$, Fig. 9 presents t-SNE [70] visualization of the class token on the Animal3D test set. It is evident that $\mathcal{L}_{\text{con}}$ brings features of animals from the same family closer, thereby enhancing the model’s ability to distinguish between the shapes of animals from different families. Fig. 10 further demonstrates that $\mathcal{L}_{\text{con}}$ enables the model to better distinguish between different animal families. Specifically, when trained without $\mathcal{L}_{\text{con}}$, the model incorrectly predicts a zebra, a bear, and an antelope as a dog, a zebra, and a cow, respectively.
@@ V-E3 Ablations on AniMer+
TABLE XII: Effect of MoE. PAJ: PA-MPJPE. PAV: PA-MPVPE. P@0.1: PCK@0.1. AniMer-AM: training without MoE. AniMer+: training with MoE.
@@
To demonstrate that AniMer+ is more effective in simultaneously handling animals with anatomical differences, we compare the results of AniMer+ training on both mammals and birds with the results of AniMer training on both mammals and birds. We use ViT (AniMer-AM) and ViT-MoE (AniMer+) as the backbones respectively. Note that AniMer-AM and AniMer+ have the same number of parameters. The results presented in Table XII demonstrate that by incorporating an MoE design, AniMer+ is more effective at simultaneously managing animals with significant gaps in physiological structure.
@@ VI More results and analyses
Comparison Between CtrlAni3D and CG synthetic. To further emphasize the superiority of our diffusion based gerneration pipeline over traditional CG synthesization, we follow the experiments performed by Animal3D [18], as tabulated in Table XIII. Specifically, “HMR” means directly training HMR on Animal3D. “HMR-Synthetic” pretrains HMR with a CG-based synthetic dataset generated by textured SMAL [42] for 100 epochs before training on Animal3D. Similarly, “HMR-CtrlAni3D” pretrains HMR on CtrlAni3D for 100 epochs. It is evident that CtrlAni3D provides better pretraining than traditional CG based synthetic data.
TABLE XIII: CtrlAni3D versus CG synthetic. The values of “HMR-Synthetic” are borrowed from paper  [18].
@@
Domain Gap Between CtrlAni3D and Real-world Data. Although CtrlAni3D can improve model performance on real-world data, there still exists certain domain gap. Table XIV indicates certain domain gap between Animal3D and CtrlAni3D. However, the comparable results between AniMer-M (A3D) and AniMer-M (C3D) on the Animal Kingdom dataset indicate a similar generalizability on in-the-wild data between Animal3D and CtrlAni3D. This is why we aggregate many different datasets for full training. The performance gains of PA-MPJPE and PA-MPVPE shown in Table VIII further validate the effectiveness of CtrlAni3D in assisting model generalization.
TABLE XIV: The generalizability of CtrlAni3D. A3D means training AniMer-M only on Animal3D, C3D means training AniMer-M only on CtrlAni3D.
@@
PCK@HTH $\uparrow$ AUC $\uparrow$ PCK@HTH $\uparrow$ AUC $\uparrow$ PCK@HTH $\uparrow$ AUC $\uparrow$
@@
Comparison Between CtrlAVES3D and CUB17. Table XV shows the results of training solely on CUB17 and training solely on CtrlAVES3D. Our CtrlAVES3D achieves better results than CUB17 plausibly because CtrlAVES3D encompasses 3D annotation and diverse scenes compared to CUB17.
TABLE XV: CUB17 versus CtrlAVES3D. CUB17 means training only on CUB17, CA3D means training only on CtrlAVES3D.
@@
PCK@HTH $\uparrow$ AUC $\uparrow$ PCK@HTH $\uparrow$ AUC $\uparrow$ PCK@HTH $\uparrow$ AUC $\uparrow$
@@
[…]
TABLE XVI: Comparison between classifier loss and contrastive loss on real-world datasets (Animal3D, Cow Bird and CUB). PAJ: PA-MPJPE, PAV: PA-MPVPE, P@0.1: PCK@0.1.
@@ VIII Conclusion
Summary. In this paper, we first introduce AniMer, a simple yet effective approach for precisely estimating animal pose and shape. Then we extend AniMer to AniMer+, which unifies the mesh recovery of two distinct animal taxa (mammalia and aves) for the first time. The key to the success of AniMer and AniMer+ is a large capacity Transformer backbone together with an aggregated large-scale dataset. For the aggregated dataset, we propose a novel synthetic dataset generation pipeline to generate two novel synthetic general datasets, CtrlAni3D for quadrupeds and CtrlAVES3D for birds, which are rendered by prompting a controllable text-to-image generation model ControlNet. Benefiting from the model design, the synthetic datasets, and our family aware design philosophy, AniMer and AniMer+ outperform previous methods not only on 3D mammalian and avian datasets included for training, but also on out-of-domain 2D real-world mammalian and avian datasets. Moreover, the MoE design also enables AniMer+ to improve the performance across anatomically diverse taxa within a single model. The principles behind both AniMer and AniMer+ shall inspire the mesh recovery tasks of the whole animal kingdom and enable several downstream applications such as avatar creation and behavioral analysis.
Limitations and Future Work. Despite the demonstrated robustness, our work highlights several avenues for future improvement. Our analysis identifies three main limitations: (1) Model Expressivity: The AVES parametric model, while effective for many poses, lacks the expressiveness to accurately represent complex avian articulations, such as those during flight. (2) Challenging Scenarios: Motion blur and severe occlusions would harm the reconstructed results. (3) Data Generation Fidelity: ControlNet’s species‑specific performance remains uneven, occasionally producing samples that do not correspond to the intended prompt. In future work, we will address these shortcomings by integrating more expressive animal parametric models, adopting spatio‑temporal reconstruction methods, and enhancing ControlNet’s understanding of diverse species. Building upon our current multi-species success, we also aim to expand the model’s applicability to an even broader and more diverse range of animals.
Acknowledgements. This study was supported by the National Key Research and Development Program of China (2023YFC2415400); the National Natural Science Foundation of China (T2422012, 62071210, 62125107); the Guangdong Basic and Applied Basic Research (2024B1515020088); the Shenzhen Science and Technology Program (RCYX20210609103056042); the High Level of Special Funds (G030230001, G03034K003); the Shuimu Tsinghua Scholar Program (2024SM324).

### v2

@@ V-B Qualitative Results of AniMer+
Figure 6: Qualitative comparisons on the Animal3D, CtrlAni3D, and Animal Kingdom datasets. We compare our results with HMR [kanazawa2018end], WLDO [biggs2020wldo], AniMer-a (ResNet152 backbone), AniMer-b (no pretraining) and HMR2.0 [goel2023humans].
Figure 7: Qualitative comparisons on the CUB, CtrlAVES3D and Cow Bird datasets. We compare our results with AVES [wang2021birds] and HMR2.0 [goel2023humans].
@@ V-C1 Mammalian Experiments
Following the practice of the Animal3D dataset [xu2023animal3d], we utilize HMR [kanazawa2018end] and WLDO [biggs2020wldo] as baselines. Also, we compare with HMR2.0 [goel2023humans] to highlight our special design choices. To be fair, all compared methods are trained on our full training dataset with the same losses and the same two-stage training strategy.
@@
In addition, we compare our method with a typical template-free method 3D Fauna [li2024learning] on the Animal3D dataset. The PCK@0.1 results for 3D Fauna are sourced from the original paper. Note that, the PCK@0.1 of 3D Fauna uses square root of the 2D bounding box area as a normalization factor. For fair comparison, we use the same threshold for PCK@0.1 calculation in this experiment. Besides, due to the inconsistency between the mesh topology of the 3D Fauna output and SMAL, we utilize PA-CD as the 3D evaluation metric instead of PA-MPJPE and PA-MPVPE. The quantitative results are displayed in Table V. AniMer-M and AniMer+ demonstrate comparable performance to 3D Fauna on PCK@0.1. For PA-CD, AniMer-M and AniMer+ show improvements of $68.2\%$ and $67.4\%$, respectively, compared to 3D Fauna.
@@ V-C2 Avian Experiments
We conduct both quantitative and qualitative comparisons against AVES [wang2021birds] and HMR2.0 [goel2023humans]. Unlike mammals, birds exhibit a more limited range of poses and shapes, so the model converges faster. Therefore, we train AniMer-A in a single stage. The quantitative results in Table VI show that AniMer-A and AniMer+ outperform all baselines across all datasets. Notably, our method achieves the best results on the Cow Bird dataset despite its large number of blurred images and the fact that this dataset is not seen during training—underscoring the robustness of our method. While AniMer+ and AniMer-A perform comparably on our synthetic CtrlAVES3D dataset, the specialized AniMer-A achieves higher accuracy on the real-world CUB and Cow Bird datasets, as its entire network capacity is dedicated to avian reconstruction.
@@ V-D2 Ablations on CtrlAVES3D
Similar to CtrlAni3D, we conduct an experiment to demonstrate the effectiveness of incorporating CtrlAVES3D training for out-of-domain datasets. The results are shown in Table IX. Since the Cow Bird dataset includes a variety of viewpoints, and CtrlAVES3D encompasses diverse perspectives and 3D annotations, models trained with CtrlAVES3D outperform those trained without it. When compared to training solely on CUB or CtrlAVES3D, training on both CUB and CtrlAVES3D results in improvements of $18.7\%$ and $4.6\%$ in AUC, respectively.
V-D3 Ablations on Filtering Strategy
The filtering strategy directly contributes to the quality of our synthetic datasets, which in turn significantly influences the model’s performance and generalization. We compare training with failure cases (“w/ failure cases” in Table X) and training without failure cases (“w/o failure cases” in Table X) on single animal family Equidae (horse and zebra in our case).The Equidae subset of CtrlAni3D contains totally 3138 successful cases for horse and zebra. We further purposely generate 1326 failure cases. Both models are trained for 500 epochs to converge. The substantially improved results in Table X demonstrate the effectiveness of the filtering strategy.
TABLE X: The impact of training with failure cases. The results are obtained by evaluating on the Animal3D testset.
Horse Zebra
w/o failure cases w/ failure cases w/o failure cases w/ failure cases
$downarrow PA-MPJPE\downarrow PA-MPVPE\downarrow$ $downarrow PA-MPJPE\downarrow PA-MPVPE\downarrow$
107.9 118.3 141.8 141.9 117.0 119.6 145.6 165.1
V-D4 Experiments on Challenging Poses
Synthetic pose and shape datasets may encounter a sim-to-real gap, especially for pose diversity. To test the efficacy of our datasets for rare and challenging poses, we manually select 310 images with challenging poses (e.g., lying down) from the Animal Kingdom dataset to validate Animer-M models trained with or without CtrlAni3D (“w/ CtrlAni3D” vs. “w/o CtrlAni3D” in Table XI). Note that, we choose the Animal Kingdom dataset because their labels do not relate to SMAL, and therefore the challenging poses in this dataset are more likely to stay outside the sampled pose space of SMAL. The results in Table XI prove that CtrlAni3D helps reconstruct rare poses better, even if sampling is not specifically conducted for challenging postures.
TABLE XI: The impact of CtrlAni3D for challenging poses. The results are obtained by evaluating on a subset of 310 challenging images from the Animal Kingdom dataset.
@@ V-D4 Experiments on Challenging Poses
w/o CtrlAni3D 80.6 29.6 48.0
w/ CtrlAni3D 81.0 30.4 50.2
V-D5 Experiments on Dataset Size of CtrlAni3D
The current training sizes of CtrlAni3D (8,277 images) and CtrlAVES3D (6,464 images) are primarily determined by trading off computational resources, generation time, and observed performance during preliminary experiments. To investigate the behavior of the model’s performance as the synthetic data size scales, we progressively increase the size of CtrlAni3D from 2.5K to 10K for training with 500 epochs and test on the Animal3D test set. Please note for this scalability analysis, only CtrlAni3D is utilized as the training data. Table XII clearly demonstrates a positive correlation between synthetic dataset size and model performance. However, when the dataset size exceeds 5K, the performance gain slows down.
TABLE XII: Dataset scalability analysis on CtrlAni3D. The results are obtained by training on CtrlAni3D and testing on the Animal3D testset.
[…]
2.5K 79.5 113.5
5K 82.3 107.1
7.5K 82.8 106.9
10K 83.5 106.7
@@ V-E1 Ablations on Different Encoder and Decoder
To emphasize the significance of the ViT encoder, backbone pretraining and the Transformer decoder, we substitute the ViT encoder with a ResNet-152 encoder, discard ViT pretraining and replace the Transformer decoder with an MLP decoder, respectively. The results are presented in Table XIII. Compared with traditional methods, which use CNN as a backbone and an MLP head as a decoder, our ViT encoder and Transformer decoder improve the performance of the model’s output mesh and image alignment. In addition, a backbone with random initialization but no pretraining leads to poor performance and is hard to converge.
TABLE XIII: Ablations on different encoder and decoder. AniMer-a: use ResNet-152 as the encoder. AniMer-b: discard ViT pretraining. AniMer-c: use MLP as the decoder. PAJ: PA-MPJPE, PAV: PA-MPVPE, P@0.1: PCK@0.1.
@@
PAJ $\downarrow$ PAV $\downarrow$ PAJ $\downarrow$ PAV $\downarrow$ AUC $\uparrow$ P@0.1 $\uparrow$
@@
V-E2 Ablations on Animal Family Supervised Contrastive Learning
TABLE XIV: Effect of animal family supervised contrastive learning. PAJ: PA-MPJPE. PAV: PA-MPVPE. P@0.1: PCK@0.1.
@@ V-E2 Ablations on Animal Family Supervised Contrastive Learning
We conduct experiments to demonstrate the effect of animal family supervised contrastive learning. Table XIV presents the quantitative results on both mammalian and avain datasets. Clearly, AniMer-M trained with $\mathcal{L}_{\text{con}}$ outperforms the model that does not utilize $\mathcal{L}_{\text{con}}$. The improved PA-MPVPE and PA-MPJPE metrics indicate that $\mathcal{L}_{\text{con}}$ enhances the model’s capability in shape and pose estimation.
@@
In addition, to provide more intuitive understanding of the effect of $\mathcal{L}_{\text{con}}$, Fig. 9 presents t-SNE [van2008visualizing] visualization of the class token on the Animal3D test set. It is evident that $\mathcal{L}_{\text{con}}$ brings features of animals from the same family closer, thereby enhancing the model’s ability to distinguish between the shapes of animals from different families. Fig. 10 further demonstrates that $\mathcal{L}_{\text{con}}$ enables the model to better distinguish between different animal families. Specifically, when trained without $\mathcal{L}_{\text{con}}$, the model incorrectly predicts a zebra, a bear, and an antelope as a dog, a zebra, and a cow, respectively.
@@ V-E3 Ablations on AniMer+
TABLE XV: Effect of MoE. PAJ: PA-MPJPE. PAV: PA-MPVPE. P@0.1: PCK@0.1. AniMer-AM: training without MoE. AniMer+: training with MoE.
@@
To demonstrate that AniMer+ is more effective in simultaneously handling animals with anatomical differences, we compare the results of AniMer+ training on both mammals and birds with the results of AniMer training on both mammals and birds. We use ViT (AniMer-AM) and ViT-MoE (AniMer+) as the backbones respectively. Note that AniMer-AM and AniMer+ have the same number of activated parameters during a single pass of inference. The results presented in Table XV demonstrate that by incorporating an MoE design, AniMer+ is more effective at simultaneously managing animals with significant gaps in physiological structure.
V-E4 Ablations on the Feature Dimensions of MoE
We choose a feature dimension of 320 for taxa-specific layer and 960 for taxa-shared layer (1280 in total), resulting in a 1:3 specific-to-shared dimension ratio for MoE. We conduct an ablation study in Table XVI to compare the performance of different specific-to-shared dimension ratios within the fixed 1280 full dimension. For 1:1 ratio, both specific and shared dimensions are 640. For 1:4 ratio, the specific dimension is 256 while the shared dimension is 1024. Note that we do not test 1:2 and 1:5 because 1280 can not be divided by 3 nor 6 evenly. From the results tabulated in that table, the 1:3 ratio (320 specific and 960 shared) consistently achieves the best performance across all datasets.
TABLE XVI: Ablation study on feature dimension ratios of taxa-specific and taxa-shared layers. PAJ: PA-MPJPE. PAV: PA-MPVPE. P@0.1: PCK@0.1.
Ratio Animal3D CtrlAni3D Animal Kingdom
[…]
1:1 82.8 83.0 47.1 50.0 81.0 32.8
1:4 79.0 83.5 48.0 51.4 81.5 33.0
1:3 78.0 82.7 44.1 47.6 82.6 33.1
Ratio CtrlAVES3D CUB Cow Bird
[…]
1:1 67.4 71.2 88.5 56.3 57.5 16.0
1:4 66.0 70.9 88.1 55.0 59.0 15.5
1:3 65.6 70.9 89.5 59.4 59.3 16.4
V-E5 Ablations on Loss Weights
The weights for different loss terms are chosen empirically with the general goal of balancing their contributions to the total loss. Given the large number of hyperparameters, a comprehensive grid search for all loss weights is not feasible. Therefore, we only perform ablation study for the weight of our proposed animal family supervised contrastive loss $\mathcal{L}_{con}$, as it is a key component of our method. Experiments are conducted on the Animal Kingdom dataset. As shown in Table XVII, we test three different values for $\lambda_{con}$.
TABLE XVII: Ablation study on the weight of the contrastive loss, evaluated on the OOD Animal Kingdom dataset.
[…]
0.01 82.6 33.8 53.4
0.001 83.0 36.3 54.2
0.0005 82.9 34.9 54.7
While the 0.001 weight achieves the best results on AUC and PCK@0.1 metrics, the 0.0005 weight yields the best performance on the PCK@0.15 metric and is competitive on other metrics. Therefore, for all other experiments, we set $\lambda_{con}=0.0005$.
V-E6 Ablations on Two-Stage Training
As stated before, we employ two-stage training for mammalian models and single-stage training for avian models. The difference stems from the composition of the datasets. Mammalian datasets are large-scaled and have an imbalanced 2D-to-3D data ratio, while avian datasets are smaller-scaled and more balanced. We conduct ablation studies on the stage choices for both mammalian and avian models in Table XVIII.
For the mammalian experiments, the results clearly show that the two-stage training strategy consistently and significantly outperforms the single-stage approach across all evaluation datasets (Animal3D, CtrlAni3D, and Animal Kingdom). This finding holds for both 3D metrics (e.g., PA-MPJPE on Animal3D drops from 85.9 to 80.5) and 2D metrics (e.g., AUC on Animal Kingdom increases from 80.6 to 82.9). These results validate our intuition that a two-stage process is beneficial for handling the large size and the imbalanced 2D-to-3D data ratio in the mammalian datasets.
For the avian experiments, the results demonstrates that single-stage training yields superior performance for birds. On the CUB dataset, for instance, single-stage training achieves a much higher PCK@0.1 (74.8 vs. 58.2) and a better AUC (91.9 vs. 89.4). This supports our decision to use a simpler, single-stage process for the smaller and more balanced avian datasets.
TABLE XVIII: Ablation study on the number of training stages. PAJ: PA-MPJPE. PAV: PA-MPVPE. P@0.1: PCK@0.1.
Mammalian
Experiment
Animal3D CtrlAni3D Animal Kingdom
[…]
Single-stage 85.9 90.4 59.5 64.2 80.6 28.6
Two-stage 80.5 85.7 44.1 47.6 82.9 34.9
Avain
Experiment
CtrlAVES3D CUB Cow Bird
[…]
Single-stage 65.5 90.2 91.9 74.8 65.7 19.4
Two-stage 71.3 90.3 89.4 58.2 63.0 20.1
@@ VI More results and analyses
Comparison Between CtrlAni3D and CG synthetic. To further emphasize the superiority of our diffusion based gerneration pipeline over traditional CG synthesization, we follow the experiments performed by Animal3D [xu2023animal3d], as tabulated in Table XIX. Specifically, “HMR” means directly training HMR on Animal3D. “HMR-Synthetic” pretrains HMR with a CG-based synthetic dataset generated by textured SMAL [zuffi2018lions] for 100 epochs before training on Animal3D. Similarly, “HMR-CtrlAni3D” pretrains HMR on CtrlAni3D for 100 epochs. It is evident that CtrlAni3D provides better pretraining than traditional CG based synthetic data.
TABLE XIX: CtrlAni3D versus CG synthetic. The values of “HMR-Synthetic” are borrowed from paper  [xu2023animal3d].
@@
Domain Gap Between CtrlAni3D and Real-world Data. Although CtrlAni3D can improve model performance on real-world data, there still exists certain domain gap. Table XX indicates certain domain gap between Animal3D and CtrlAni3D. However, the comparable results between AniMer-M (A3D) and AniMer-M (C3D) on the Animal Kingdom dataset indicate a similar generalizability on in-the-wild data between Animal3D and CtrlAni3D. This is why we aggregate many different datasets for full training. The performance gains of PA-MPJPE and PA-MPVPE shown in Table VIII further validate the effectiveness of CtrlAni3D in assisting model generalization.
TABLE $X$: The generalizability of CtrlAni3D. A3D means training AniMer-M only on Animal3D, C3D means training AniMer-M only on CtrlAni3D. P@H: PCK@HTH.
@@
P@H $\uparrow$ AUC $\uparrow$ P@H $\uparrow$ AUC $\uparrow$ P@H $\uparrow$ AUC $\uparrow$
@@
Comparison Between CtrlAVES3D and CUB17. Table XXI shows the results of training solely on CUB17 and training solely on CtrlAVES3D. Our CtrlAVES3D achieves better results than CUB17 plausibly because CtrlAVES3D encompasses 3D annotation and diverse scenes compared to CUB17.
TABLE XXI: CUB17 versus CtrlAVES3D. CUB17 means training only on CUB17, CA3D means training only on CtrlAVES3D. P@H: PCK@HTH.
@@
P@H $\uparrow$ AUC $\uparrow$ P@H $\uparrow$ AUC $\uparrow$ P@H $\uparrow$ AUC $\uparrow$
@@
More Discussion About Contrastive Learning. To further demonstrate the effectiveness of our family-aware contrastive learning scheme, we experiment to compare $\mathcal{L}_{\text{con}}$ with $\mathcal{L}_{\text{cls}}$, which simply replaces supervise contrastive loss with cross-entropy loss. Specifically, the contrastive loss directly impacts the feature tokens $\mathbf{F}$, which in turn indirectly impacts the feature vectors $\bm{f}$ and aligns features to model the global structure, capturing family differences. This ensures that the final output shape aligns more closely with the category of the input image. Compared with contrastive learning, $\mathcal{L}_{\text{cls}}$ (“w $\mathcal{L}_{\text{cls}}$” in Table XXII) focuses solely on optimizing classification accuracy, which may not necessarily improve geometric parameter regression. This is why in Table XXII we observe that training with $\mathcal{L}_{\text{cls}}$ may slightly deteriorate pose and shape estimation. Moreover, contrastive learning facilitates a more compact intra-class distribution and a more separable inter-class distribution in the feature space [khosla2020supervised], thereby enhancing the model’s capability for few-shot learning. In Table XXII, we report the results of AniMer-M and AniMer-A on various specific animal species. It is clearly shown that $\mathcal{L}_{\text{con}}$ can improve performance for animals with limited training samples (e.g., boars constitute less than $1\%$ of the training set).
TABLE XXII: Comparison between classifier loss and contrastive loss on real-world datasets (Animal3D, Cow Bird and CUB). PAJ: PA-MPJPE, PAV: PA-MPVPE, P@0.1: PCK@0.1.
@@ VIII Conclusion
Summary. In this paper, we first introduce AniMer, a simple yet effective approach for precisely estimating animal pose and shape. Then we extend AniMer to AniMer+, which unifies the mesh recovery of two distinct animal taxa (mammalia and aves) for the first time. The key to the success of AniMer and AniMer+ is a large capacity Transformer backbone together with an aggregated large-scale dataset. For the aggregated dataset, we propose a novel synthetic dataset generation pipeline to generate two novel synthetic general datasets, CtrlAni3D for quadrupeds and CtrlAVES3D for birds, which are rendered by prompting a controllable text-to-image generation model ControlNet. Benefiting from the model design, the synthetic datasets, and our family aware design philosophy, AniMer and AniMer+ outperform previous methods not only on 3D mammalian and avian datasets included for training, but also on out-of-domain 2D real-world mammalian and avian datasets. Moreover, the MoE design also enables AniMer+ to improve the performance across anatomically diverse taxa within a single model. Our work serves as a blueprint that can readily incorporate new parametric models as they become available. The principles behind both AniMer and AniMer+ shall inspire the mesh recovery tasks of the whole animal kingdom and enable several downstream applications such as avatar creation and behavioral analysis.
Limitations and Future Work. Despite the demonstrated robustness, our work highlights several avenues for future improvement. Our analysis identifies three main limitations: (1) Model Expressivity: The AVES parametric model, while effective for many bird poses, lacks the expressiveness to accurately represent complex avian articulations, such as those during flight. For mammalia, the pose space of the SMAL model is based on a limited set of species. For other animal species like primates or rodents, there is a lack of effective parametric models. (2) Challenging Scenarios: Motion blur and severe occlusions would harm the reconstructed results. (3) Data Generation Fidelity: ControlNet’s species‑specific performance remains uneven, occasionally producing samples that do not correspond to the intended prompt.
In future work, we will address these shortcomings by developing more expressive animal parametric models, incorporating more varied motion capture data from other species, adopting spatio‑temporal reconstruction methods, and enhancing ControlNet’s understanding of diverse species. Building upon our current multi-species success, we also aim to expand the model’s applicability to an even broader and more diverse range of animals.
Acknowledgements. This work was done at the Department of Electronic and Electrical Engineering, Southern University of Science and Technology, Shenzhen, China. This study was supported by the National Key Research and Development Program of China (2023YFC2415400); the National Natural Science Foundation of China (T2422012, 62071210, 62125107); the Guangdong Basic and Applied Basic Research (2024B1515020088); the Shenzhen Science and Technology Program (RCYX20210609103056042); the High Level of Special Funds (G030230001, G03034K003); the Shuimu Tsinghua Scholar Program (2024SM324).